# Bayesian Characterization and Identifiability of Multilevel Quantum Systems

### v1

@@ Abstract
Robust control of a quantum system is essential to utilize the current noisy quantum hardware to their full potential, such as quantum algorithms. To achieve such a goal, systematic search for an optimal control for any given experiment is essential. Design of optimal control pulses require accurate numerical models, and therefore, accurate characterization of the system parameters. We present an online, Bayesian approach for quantum characterization of qutrit systems which automatically and systematically identifies the optimal experiments that provide maximum information on the system parameters, thereby greatly reducing the number of experiments that need to be performed on the quantum testbed. Unlike most characterization protocols that provide point-estimates of the parameters, the proposed approach is able to estimate their probability distribution. The applicability of the Bayesian experimental design technique was demonstrated on test problems where each experiment was defined by a parameterized control pulse. In addition to this, we also presented an approach for adaptive pulse parameterization which is robust under uncertainties in transition frequencies and coherence times, and shot noise, despite being initialized with wide uninformative priors. Furthermore, we provide a mathematical proof of the theoretical identifiability of the model parameters and present conditions on the quantum state under which the parameters are identifiable. The proof and conditions for identifiability are presented for both closed and open quantum systems using the Schroedinger equation and the Lindblad master equation respectively.
@@ 1 Introduction
The BExD approach has been used to identify parameters in electro-chemical models of Lithium-Ion cells [12], pharmacokinetic model [13], Thomson scattering model for nuclear fusion [14] and electrical impedance tomography [15]. The probabilistic nature of quantum mechanics naturally lends itself to Bayesian characterization. The BExD technique has been used to estimate the resonant frequency of a single two-level closed system under time-independent Hamiltonian by minimizing the Bayes’ risk [16]. Granade et al.[17] demonstrated the BExD approach to infer the transition frequency and $T_{2}$ dephasing term on a canonical, two-level system, and built the open-source framework QInfer [18]. McMichael et al.[19] utilized BExD to determine optimal controls in Ramsey interferometry to estimate precession frequency and $T_{2}^{\star}$ decoherence time, and demonstrated a five-factor speed up over random sampling methods. Wang et al.[20] demonstrated the BExD framework for quantum sensing of the environment around a Nitrogen-Vacancy (NV) center in diamond and reported a 90% speed up over their heuristic method in estimating the nearby nuclear spins and oscillating magnetic field. Hincks et al.[21] estimated five-parameters in their Hamiltonian and Lindblad model for NV in diamonds with Ramsey and Rabi style experiments where they showed the BExD approach results in a posterior variance that is two orders of magnitude smaller than those obtained using the heuristic methods. Wang et al.[22] performed characterization of an NV center in diamonds by estimating the Rabi frequency and updating their Hamiltonian model to include missing interactions (chirping). Both Wang et al.[22] and Ferrie et al.[16] demonstrated exponential convergence of the BExD (when minimizing posterior variance) up to a saturation point, beyond which the ansatz is no longer able to model the missing interactions. Although paramount, most studies for characterization of the quantum system often do not consider the identifiability of the parameters under a given experimental setup. Wang et al.[23] demonstrated that the identifiability of closed quantum systems with a single qubit in density matrix formalism (Liouville-von Neumann equation). Zhang et al.[24] investigated the identifiability of both closed and open, dephasing qubit systems and demonstrated the identifiability can be greatly increased by performing appropriate experiments that can be parameterized by control pulses driving the quantum system.
@@
However, the identifiability of multi-level closed and open quantum systems has yet to be explored. Hence, this work demonstrates the identifiability of parameters in the Schroedinger and Lindblad models of multi-level closed and open (with energy decay and dephasing) systems, respectively, and the sufficient conditions on quantum state for the parameters to be identifiable under a set of measurements. Furthermore, it demonstrates the invariance of the conditions to unitary transformations, hence their applicability to a different set of measurements. The BExD approach is then used to identify optimal experiments. Identifying these optimal experimental parameters amounts to maximizing functions of the experiment controls, parameters, and prior that quantifies how informative the experiment would be if performed. These utility functions will be discussed in the following sections. Although demonstrated on a single, qubit problems, its performance for characterizing quantum systems with higher energy levels has yet to be thoroughly investigated. This work aims to address this scarcity and apply the BExD technique for characterizing multi-level quantum systems. We present an adaptive pulse parameterization to accurately estimate parameters that govern processes operating on different time scales.
@@ 2 Quantum Mechanical Model
where the rotation frequency in the RWA is taken to be the drive frequency $\omega_{d}$, and $\Delta:=\omega-\omega_{d}$ is the detuning between the qubit 0-1 transition and the drive frequencies. The decoherence is introduced through the Lindblad jump operators for decay ($L_{1}:=\frac{1}{\sqrt{T_{1}}}a$) and dephasing ($L_{2}:=\frac{1}{\sqrt{T_{2}}}a^{\dagger}a$) where $T_{1}$ and $T_{2}$ are the decay and dephasing times, respectively. The Lindbladian is then written as
@@ 3 Identifiability of Quantum Mechanical Models
where $m$ is the smallest positive integer such that the resulting system is solvable for the parameters. This minimum order of expansion is not known a priori but as will be shown, first order and second order expansions are sufficient for closed and open quantum systems, respectively.
@@ 3.1 Model Identifiability: Schroedinger Equation
Consider a closed quantum system modeled by the Schroedinger equation $\dot{\left|\psi\right\rangle}=-iH\left|\psi\right\rangle$, driven by a known pulse $f(t)$ with the probability of measuring state $\left|i\right\rangle$ by $y_{i}(t;\mathrm{\boldsymbol{\theta}}):=\mathrm{Pr}{}(\left|i\right\rangle,t;\mathrm{\boldsymbol{\theta}})=\left\langle\mathrm{\boldsymbol{\psi}}(t;\mathrm{\boldsymbol{\theta}})\right|M_{i}\left|\mathrm{\boldsymbol{\psi}}(t;\mathrm{\boldsymbol{\theta}})\right\rangle$. We expand the measurement about $t_{0}=0$ with $\mathcal{F}:=f(t_{0})$, which yields the first three expansion coefficients
@@
$\underbrace{\begin{pmatrix}\mathrm{Re}(\psi_{0}\overline{\psi_{1}})&0\\ \mathrm{Re}(\psi_{0}\overline{\psi_{1}})-\sqrt{2}\mathrm{Re}(\psi_{1}\overline{\psi_{2}})&\sqrt{2}\mathrm{Re}(\psi_{1}\overline{\psi_{2}})\\ \mathrm{Re}(\psi_{1}\overline{\psi_{2}})&-\mathrm{Re}(\psi_{1}\overline{\psi_{2}})\end{pmatrix}}_{A}\begin{Bmatrix}\omega\\ \chi\end{Bmatrix}=\mathcal{S}$ (13)
Notice that the rows are linearly dependent. Since the drive $\mathcal{F}$ is known (or can be identified as indicated by the first order terms), the sufficient conditions to identify $\omega$ and $\chi$ are $\mathrm{Re}(\psi_{0}\overline{\psi_{1}})\neq 0$ and $\mathrm{Re}(\psi_{1}\overline{\psi_{2}})\neq 0$, respectively. Furthermore, the sufficient conditions for identifiability of $\chi$ requires that the conditions for $\omega$ also be satisfied. Hence, the parameters are globally and uniquely identifiable and the model is globally structrually identifiable, when the control pulse is able to prepare a state that satisfies these conditions.
@@ 3.2 Model Identifiability: Lindblad Master Equation
$\underbrace{\begin{pmatrix}2\mathrm{Im}(\rho_{10})&\rho_{11}\\ \sqrt{2}\mathrm{Im}(\rho_{21})&\rho_{22}\end{pmatrix}}_{A}\begin{Bmatrix}\mathcal{F}\\ \tau_{1}\end{Bmatrix}=\mathcal{S}_{1}$ (17)
Hence, if $A$ is non-singular, the first order terms prove global and unique identifiability of $\mathcal{F}$ and $\tau_{1}$. The second order terms are given as
@@
$\underbrace{\begin{pmatrix}2\mathcal{F}\mathrm{Re}(\rho_{10})&0&-\mathcal{F}\mathrm{Im}(\rho_{10})\\ 2\mathcal{F}\left(\mathrm{Re}(\rho_{10})-\sqrt{2}\mathrm{Re}(\rho_{21})\right)&2\sqrt{2}\mathcal{F}\mathrm{Re}(\rho_{21})&A_{12}\\ 2\sqrt{2}\mathcal{F}\mathrm{Re}(\rho_{21})&-2\sqrt{2}\mathcal{F}\mathrm{Re}(\rho_{21})&A_{22}\end{pmatrix}}_{A}\begin{Bmatrix}\omega\\ \chi\\ \tau_{2}\end{Bmatrix}=\mathcal{S}_{2}$ (19)
for the remaining unknown parameters where the remaining terms involving the known or identifiable parameters ($\mathcal{F}$ and $\tau_{1}$) are grouped in $\mathcal{S}_{2}\in\mathbb{R}^{3}$. These second order terms show that $\omega,~{}\chi,~{}\tau_{2}$ are globally identifiable if $A$ is non-singular. Hence, the control function should be able to drive the system to such a state. Similarly, the identifiability of the density matrix formalism for closed system (i.e., Liouville-von Neumann equation) can be determined by fixing $\tau_{1}=\tau_{2}=0$. In this case, the first order conditions only show identifiability of $\mathcal{F}$, like in the case of state vector formalism (i.e., Schroedinger equation §3.1), and a second order expansion is needed to determine identifiability of the remaining parameters $\omega$, $\chi$. Note that these identifiability conditions in both density matrix and state vector formalism are equivalent. Letting $\rho(t_{0})=\left|\mathrm{\boldsymbol{\psi}}(t_{0})\right\rangle\left\langle\mathrm{\boldsymbol{\psi}}(t_{0})\right|$, we obtain the condition for $\mathrm{Im}(\rho_{ij})=\mathrm{Im}(\psi_{i}\overline{\psi_{j}})\neq 0$ for $i\neq j$ for parameter $\mathcal{F}$ , $\mathrm{Re}(\rho_{01})=\mathrm{Re}(\psi_{0}\overline{\psi_{1}})\neq 0$ for parameter $\omega$ and $\mathrm{Re}(\rho_{21})=\mathrm{Re}(\psi_{2}\overline{\psi_{1}})\neq 0$ for parameter $\chi$. We should mention that these conditions also hold for other measurement basis $\widetilde{M}$ by applying an appropriate unitary transformation $U$ such that $\widetilde{M}_{i}=UM_{i}U^{\dagger}$ since $y_{i}(t)=\mathrm{Tr}\left(UM_{i}U^{\dagger}\rho(t)\right)=\mathrm{Tr}\left(M_{i}U^{\dagger}\rho(t)U\right)=\mathrm{Tr}\left(M_{i}\widetilde{\rho}(t)\right)$ where the conditions on the quantum state now apply in the transformed basis. These conditions show that the model parameters are identifiable by measuring the population in a single basis without needing to estimate the full density matrix via state or process tomography, which can be expensive.
@@ 4 Bayesian Experimental Design
where $\delta(\cdot)$ is the Dirac delta function. The search for such optimal control $\mathrm{\boldsymbol{\xi}}^{\star}$ requires the solution to an optimization problem Eq. (25) that can be high-dimensional and typically non-convex and multi-modal.
@@
For these reasons, we solve the optimization problem using a non-gradient based differential evolution algorithm. The applicability of the evolutionary method for optimization of utility function in BExD was demonstrated by Hamada et al.[36]. An experiment is then performed using the optimal control $\mathrm{\boldsymbol{\xi}}^{\star}$, and the measurements $\mathrm{\boldsymbol{y}}$ are used to construct the likelihood and update the posterior distribution at the $n^{th}$ epoch/iteration as
@@
$\propto\mathrm{Pr}{}(\mathrm{\boldsymbol{y}}_{n}|\mathrm{\boldsymbol{\theta}};\mathrm{\boldsymbol{\xi}}_{n}^{\star})\prod_{i=0}^{n-2}\mathrm{Pr}{}(\mathrm{\boldsymbol{\theta}}|\mathrm{\boldsymbol{y}}_{i};\mathrm{\boldsymbol{\xi}}_{i}^{\star})$
with $\mathrm{Pr}{}(\mathrm{\boldsymbol{\theta}}|\mathrm{\boldsymbol{y}}_{0};\mathrm{\boldsymbol{\xi}}_{0}^{\star})=\mathrm{Pr}{}(\mathrm{\boldsymbol{\theta}})$. The expectation of the inferred posterior $\widehat{\mathrm{\boldsymbol{\theta}}}:=\mathbb{E}_{\mathrm{Pr}{}(\mathrm{\boldsymbol{\theta}}|\mathrm{\boldsymbol{y}})}\left[\mathrm{\boldsymbol{\theta}}\right]$ yields the unbiased estimator for the model parameters. Figure 1 presents the complete workflow for Bayesian experimental design.

### v2

# Identifiability and Characterization of Transmon Qutrits Through Bayesian Experimental Design

@@ Abstract
Robust control of a quantum system is essential to utilize the current noisy quantum hardware to their full potential, such as quantum algorithms. To achieve such a goal, systematic search for an optimal control for any given experiment is essential. Design of optimal control pulses require accurate numerical models, and therefore, accurate characterization of the system parameters. We present an online, Bayesian approach for quantum characterization of qutrit systems which automatically and systematically identifies the optimal experiments that provide maximum information on the system parameters, thereby greatly reducing the number of experiments that need to be performed on the quantum testbed. Unlike most characterization protocols that provide point-estimates of the parameters, the proposed approach is able to estimate their probability distribution. The applicability of the Bayesian experimental design technique was demonstrated on test problems where each experiment was defined by a parameterized control pulse. In addition to this, we also presented an approach for iterative pulse extension which is robust under uncertainties in transition frequencies and coherence times, and shot noise, despite being initialized with wide uninformative priors. Furthermore, we provide a mathematical proof of the theoretical identifiability of the model parameters and present conditions on the quantum state under which the parameters are identifiable. The proof and conditions for identifiability are presented for both closed and open quantum systems using the Schroedinger equation and the Lindblad master equation respectively.
@@ 1 Introduction
The BExD approach has been used to identify parameters in electro-chemical models of Lithium-Ion cells [12], pharmacokinetic model [13], Thomson scattering model for nuclear fusion [14] and electrical impedance tomography [15]. The probabilistic nature of quantum mechanics naturally lends itself to Bayesian characterization. The BExD technique has been used to estimate the resonant frequency of a single two-level closed system under time-independent Hamiltonian by minimizing the Bayes’ risk [16]. Granade et al.[17] demonstrated the BExD approach to infer the transition frequency and $T_{2}$ dephasing term on a canonical, two-level system, and built the open-source framework QInfer [18]. McMichael et al.[19] utilized BExD to determine optimal controls in Ramsey interferometry to estimate precession frequency and $T_{2}^{\star}$ decoherence time, and demonstrated a five-factor speed up over random sampling methods. Wang et al.[20] demonstrated the BExD framework for quantum sensing of the environment around a Nitrogen-Vacancy (NV) center in diamond and reported a 90% speed up over their heuristic method in estimating the nearby nuclear spins and oscillating magnetic field. Hincks et al.[21] estimated five-parameters in their Hamiltonian and Lindblad model for NV in diamonds with Ramsey and Rabi style experiments where they showed the BExD approach results in a posterior variance that is two orders of magnitude smaller than those obtained using the heuristic methods. Wang et al.[22] performed characterization of an NV center in diamonds by estimating the Rabi frequency and updating their Hamiltonian model to include missing interactions (chirping). Both Wang et al.[22] and Ferrie et al.[16] demonstrated exponential convergence of the BExD (when minimizing posterior variance) up to a saturation point, beyond which the ansatz is no longer able to model the missing interactions. Gester et al.[23] recently employed the BExD approach to calibrate two-qubit entangling operators for a closed trapped-ion systems using $1200\pm 500$ experiments in less than one minute. Although the BExD approach significantly reduced the number of experiments required for accurate estimation of the model parameters, the significant speed up is mainly due to the use of analytical solution to their Hamiltonian model. Stace et al.[24] demonstrated the BExD technique on single qubit and synthetic two qubit trapped-ion systems, where they estimated the detuning frequency, Rabi frequency and the coupling strength using optimized bang-bang pulses. They reported difficulty in estimating model parameters where the algorithm fails to converge which is attributed to the identifiability of the parameter due to the states prepared under the pulse parameterization. Although paramount, most studies for characterization of the quantum system often do not consider the identifiability of the parameters under a given experimental setup. Wang et al.[25] demonstrated that the identifiability of closed quantum systems with a single qubit in density matrix formalism (Liouville-von Neumann equation). Zhang et al.[26] investigated the identifiability of both closed and open, dephasing qubit systems and demonstrated the identifiability can be greatly increased by performing appropriate experiments that can be parameterized by control pulses driving the quantum system.
@@
However, the identifiability of multi-level closed and open quantum systems has yet to be explored. Hence, this work demonstrates the identifiability of parameters in the Schroedinger and Lindblad models of multi-level closed and open (with energy decay and dephasing) systems, respectively, and the sufficient conditions on quantum state for the parameters to be identifiable under a set of measurements. Furthermore, it demonstrates the invariance of the conditions to unitary transformations, hence their applicability to a different set of measurements. The BExD approach is then used to identify optimal experiments. Identifying these optimal experimental parameters amounts to maximizing functions of the experiment controls, parameters, and prior that quantifies how informative the experiment would be if performed. These utility functions will be discussed in the following sections. Although demonstrated on a single, qubit problems, its performance for characterizing quantum systems with higher energy levels has yet to be thoroughly investigated. This work aims to address this scarcity and apply the BExD technique for characterizing multi-level quantum systems. We present an approach for iterative pulse extension to accurately estimate parameters that govern processes operating on different time scales.
@@ 2 Quantum Mechanical Model
where the rotation frequency in the RWA is taken to be the drive frequency $\omega_{d}$, and $\Delta:=\omega-\omega_{d}$ is the detuning between the qubit 0-1 transition and the drive frequencies. This model was experimentally validated for a transmon qutrit at Quantum Device and Integration Testbed (QuDIT) at LLNL [28]. The decoherence is introduced through the Lindblad jump operators for decay ($L_{1}:=\frac{1}{\sqrt{T_{1}}}a$) and dephasing ($L_{2}:=\frac{1}{\sqrt{T_{2}}}a^{\dagger}a$) where $T_{1}$ and $T_{2}$ are the decay and dephasing times, respectively. The Lindbladian is then written as
@@ 3 Identifiability of Quantum Mechanical Models
where $m$ is the smallest positive integer such that the resulting system is solvable for the parameters. We emphasize that the order of expansion here is used only to establish identifiability conditions. This minimum order of expansion is not known a priori but as will be shown, first order and second order expansions are sufficient for closed and open quantum systems, respectively. We should mention that the identifiability conditions are derived for models in the lab-frame and, therefore, do not make any approximations such as neglecting the rapidly oscillating terms in the RWA.
@@ 3.1 Model Identifiability: Schroedinger Equation
Consider a closed quantum system modeled by the Schroedinger equation $\dot{\left|\psi\right\rangle}=-iH\left|\psi\right\rangle$, driven by a known pulse $f(t)$ with a pulse amplitude $\Omega(t)$ and the probability of measuring state $\left|i\right\rangle$ by $y_{i}(t;\mathrm{\boldsymbol{\theta}}):=\mathrm{Pr}{}(\left|i\right\rangle,t;\mathrm{\boldsymbol{\theta}})=\left\langle\mathrm{\boldsymbol{\psi}}(t;\mathrm{\boldsymbol{\theta}})\right|M_{i}\left|\mathrm{\boldsymbol{\psi}}(t;\mathrm{\boldsymbol{\theta}})\right\rangle$. The pulse amplitude $\Omega(t)$ is taken to be constant for all times. An alternative parameterization for $\Omega(t)$ can also be used, in which case, the resulting Taylor series expansion will yield additional pulse dependent terms that simply scale the coefficients $a_{i,k}$ in Eq. (8) and, therefore, do not alter the identifiability conditions. We expand the measurement about $t_{0}=0$ with $\mathcal{F}:=f(t_{0})$, which yields the first three expansion coefficients
@@
$\underbrace{\begin{pmatrix}\mathrm{Re}(\psi_{0}\overline{\psi_{1}})&0\\ \mathrm{Re}(\psi_{0}\overline{\psi_{1}})-\sqrt{2}\mathrm{Re}(\psi_{1}\overline{\psi_{2}})&\sqrt{2}\mathrm{Re}(\psi_{1}\overline{\psi_{2}})\\ \mathrm{Re}(\psi_{1}\overline{\psi_{2}})&-\mathrm{Re}(\psi_{1}\overline{\psi_{2}})\end{pmatrix}}_{A_{c}}\begin{Bmatrix}\omega\\ \chi\end{Bmatrix}=\mathcal{S}$ (13)
Notice that the rows of $A_{c}$ are linearly dependent. Since the drive $\mathcal{F}$ is known (or can be identified as indicated by the first order terms), the sufficient conditions to identify $\omega$ and $\chi$ are $\mathrm{Re}(\psi_{0}\overline{\psi_{1}})\neq 0$ and $\mathrm{Re}(\psi_{1}\overline{\psi_{2}})\neq 0$, respectively. Furthermore, the sufficient conditions for identifiability of $\chi$ requires that the conditions for $\omega$ also be satisfied. Hence, the parameters are globally and uniquely identifiable and the model is globally structrually identifiable, when the control pulse is able to prepare a state that satisfies these conditions.
@@ 3.2 Model Identifiability: Lindblad Master Equation
$\underbrace{\begin{pmatrix}2\mathrm{Im}(\rho_{10})&\rho_{11}\\ \sqrt{2}\mathrm{Im}(\rho_{21})&\rho_{22}\end{pmatrix}}_{A_{1}}\begin{Bmatrix}\mathcal{F}\\ \tau_{1}\end{Bmatrix}=\mathcal{S}_{1}$ (17)
Hence, if $A_{1}$ is non-singular, the first order terms prove global and unique identifiability of $\mathcal{F}$ and $\tau_{1}$. The second order terms are given as
@@
$\underbrace{\begin{pmatrix}2\mathcal{F}\mathrm{Re}(\rho_{10})&0&-\mathcal{F}\mathrm{Im}(\rho_{10})\\ 2\mathcal{F}\left(\mathrm{Re}(\rho_{10})-\sqrt{2}\mathrm{Re}(\rho_{21})\right)&2\sqrt{2}\mathcal{F}\mathrm{Re}(\rho_{21})&A_{12}\\ 2\sqrt{2}\mathcal{F}\mathrm{Re}(\rho_{21})&-2\sqrt{2}\mathcal{F}\mathrm{Re}(\rho_{21})&A_{22}\end{pmatrix}}_{A_{2}}\begin{Bmatrix}\omega\\ \chi\\ \tau_{2}\end{Bmatrix}=\mathcal{S}_{2}$ (19)
for the remaining unknown parameters where the remaining terms involving the known or identifiable parameters ($\mathcal{F}$ and $\tau_{1}$) are grouped in $\mathcal{S}_{2}\in\mathbb{R}^{3}$. These second order terms show that $\omega,~{}\chi,~{}\tau_{2}$ are globally identifiable if $A_{2}$ is non-singular. Hence, the control function should be able to drive the system to such a state. Similarly, the identifiability of the density matrix formalism for closed system (i.e., Liouville-von Neumann equation) can be determined by fixing $\tau_{1}=\tau_{2}=0$. In this case, the first order conditions only show identifiability of $\mathcal{F}$, like in the case of state vector formalism (i.e., Schroedinger equation §3.1), and a second order expansion is needed to determine identifiability of the remaining parameters $\omega$, $\chi$. Note that these identifiability conditions in both density matrix and state vector formalism are equivalent. Letting $\rho(t_{0})=\left|\mathrm{\boldsymbol{\psi}}(t_{0})\right\rangle\left\langle\mathrm{\boldsymbol{\psi}}(t_{0})\right|$, we obtain the condition for $\mathrm{Im}(\rho_{ij})=\mathrm{Im}(\psi_{i}\overline{\psi_{j}})\neq 0$ for $i\neq j$ for parameter $\mathcal{F}$ , $\mathrm{Re}(\rho_{01})=\mathrm{Re}(\psi_{0}\overline{\psi_{1}})\neq 0$ for parameter $\omega$ and $\mathrm{Re}(\rho_{21})=\mathrm{Re}(\psi_{2}\overline{\psi_{1}})\neq 0$ for parameter $\chi$. We should mention that these conditions also hold for other measurement basis $\widetilde{M}$ by applying an appropriate unitary transformation $U$ such that $\widetilde{M}_{i}=UM_{i}U^{\dagger}$ since $y_{i}(t)=\mathrm{Tr}\left(UM_{i}U^{\dagger}\rho(t)\right)=\mathrm{Tr}\left(M_{i}U^{\dagger}\rho(t)U\right)=\mathrm{Tr}\left(M_{i}\widetilde{\rho}(t)\right)$ where the conditions on the quantum state now apply in the transformed basis. These conditions show that the model parameters are identifiable by measuring the population in a single basis without needing to estimate the full density matrix via state or process tomography, which can be expensive.
@@ 4 Bayesian Experimental Design
where $\delta(\cdot)$ is the Dirac delta function. The search for such optimal control $\mathrm{\boldsymbol{\xi}}^{\star}$ requires the solution to an optimization problem Eq. (25) that can be high-dimensional and typically non-convex and multi-modal [39].
@@
For these reasons, we solve the optimization problem using a non-gradient based differential evolution algorithm. The applicability of the evolutionary method for optimization of utility function in BExD was demonstrated by Hamada et al.[40]. Schoneberger et al.[39] demonstrated the robustness of non-gradient based optimization algorithm for optimal experimental design. An experiment is then performed using the optimal control $\mathrm{\boldsymbol{\xi}}^{\star}$, and the measurements $\mathrm{\boldsymbol{y}}$ are used to construct the likelihood and update the posterior distribution at the $n^{th}$ epoch/iteration as
@@
$\propto\mathrm{Pr}{}(\mathrm{\boldsymbol{\theta}})\prod_{i=1}^{n}\mathrm{Pr}{}(\mathrm{\boldsymbol{y}}_{i}|\mathrm{\boldsymbol{\theta}};\mathrm{\boldsymbol{\xi}}_{i}^{\star})$
To reduce computational cost associated with evaluating the likelihood in regions of low posterior density, the Monte Carlo samples are resampled to redistribute the particles to regions of high posterior density. We make use of the resampling algorithm proposed by Liu and West [41]. The expectation of the inferred posterior $\widehat{\mathrm{\boldsymbol{\theta}}}:=\mathbb{E}_{\mathrm{Pr}{}(\mathrm{\boldsymbol{\theta}}|\mathrm{\boldsymbol{y}})}\left[\mathrm{\boldsymbol{\theta}}\right]$ yields the unbiased estimator for the model parameters. Figure 1 presents the complete workflow for Bayesian experimental design.